OpenStudy (anonymous):

i am having problems with this question, integral 4 to 8 ydy/y^2-2y-3, by expressing the integrand as a sum of partial fractions and evaluate the integrals, with the integration of rational functions by partial fractions

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{y}{y^2-2y-3}~dy}\)

OpenStudy (solomonzelman):

factor the denominator

OpenStudy (anonymous):

well i know you need to break the (y^2-2y-3) into the different things

OpenStudy (solomonzelman):

you just need to tell me (for now) how to factor the \(\large\color{slate}{\displaystyle y^2-2y-3}\).

OpenStudy (solomonzelman):

what are it's factors ?

OpenStudy (anonymous):

wait you cant can you

OpenStudy (solomonzelman):

yes, you can

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle (-3)+(1)=-2 \\ (-3) \times (1)=-3}\)

OpenStudy (anonymous):

is it (y-1)(y-1)?

OpenStudy (solomonzelman):

no it is not

OpenStudy (solomonzelman):

I gave you the hint

OpenStudy (anonymous):

yeah i see what you did, im just not thinking very well

OpenStudy (solomonzelman):

alright, when you factor it, tell me what factors you get.

OpenStudy (anonymous):

ok its (y-3)(y+1)

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{y}{(y+1)(y-3)}~dy}\)

OpenStudy (solomonzelman):

now, you need to solve the following: \(\large\color{slate}{\displaystyle\frac{y}{(y+1)(y-3)}=\frac{A}{y+1}+\frac{B}{y-3}}\)

OpenStudy (solomonzelman):

(note: this equation is true for any value of y)

OpenStudy (anonymous):

yes i have that

OpenStudy (solomonzelman):

now, multiply each term times (y+1)(y-3), and rewrite each term.

OpenStudy (anonymous):

ok when i get to that part i need help with the intrgral part from 4 to 8 \

OpenStudy (solomonzelman):

you are trying to skip steps/

OpenStudy (anonymous):

well i meant, when i finished the part before the integral

OpenStudy (solomonzelman):

lets just do that part right now, and then worry about the steps ahead

OpenStudy (solomonzelman):

when you multiply the above (the last thing I wrote in latex) each term, times (y+1)(y-3) what do you get /

OpenStudy (anonymous):

well i dont think they are righ but i have A = 3/4 and B = -3/2

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{y}{(y+1)(y-3)} = \frac{A}{y+1} +\frac{B}{y-3} }\) \(\large\color{slate}{ \displaystyle \left(\frac{y}{(y+1)(y-3)} = \frac{A}{y+1} +\frac{B}{y-3} \right)\times [~(y+1)(y-3) ~] }\) \(\large\color{slate}{ \displaystyle y=A(y-3)+B(y+1) }\)

OpenStudy (anonymous):

ok i got that now

OpenStudy (solomonzelman):

when y=-1 \(\large\color{slate}{ \displaystyle -1=A(-1-3)+B(-1+1) }\) \(\large\color{slate}{ \displaystyle -1=A(-4)+B(0) }\) \(\large\color{slate}{ \displaystyle -1=A(-4) }\) \(\large\color{slate}{ \displaystyle \frac{1}{4}=A }\) when y=3 \(\large\color{slate}{ \displaystyle y=A(y-3)+B(y+1) }\) \(\large\color{slate}{ \displaystyle 3=A(3-3)+B(3+1) }\) \(\large\color{slate}{ \displaystyle 3=A(0)+B(4) }\) \(\large\color{slate}{ \displaystyle 3=B(4) }\) \(\large\color{slate}{ \displaystyle \frac{3}{4}=B }\)

OpenStudy (anonymous):

ok that is what i am getting at wasnt using -1, used 1

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{y}{(y+1)(y-3)}~dy}\) \(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{A}{(y+1)}+\frac{B}{(y-3)}~dy}\) \(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{(1/4)}{(y+1)}+\frac{(3/4)}{(y-3)}~dy}\) then there a couple way to say it, which are all essentially same (1) \(\large\color{slate}{\displaystyle\frac{1}{4}\int\limits_{4}^{8}\left(\frac{1}{y+1}+\frac{3}{(y-3)}\right)~dy}\) (2) \(\large\color{slate}{\displaystyle\int\limits_{4}^{8}\frac{1}{4(y+1)}+\frac{3}{4(y-3)}~dy}\)

OpenStudy (solomonzelman):

then as you can probably see, the answers will be natural logarithms.

OpenStudy (anonymous):

yeah, so then just plug in 8 and 4 and you will get the answer?

OpenStudy (solomonzelman):

you will need to integrate first.

OpenStudy (anonymous):

by using u substitution?

OpenStudy (solomonzelman):

well, you can use a substitution from y+1 and y-1 in the denominators, but in the cases of y+1 and y-1, if you try to set u=y+1 and/or u=y-1, then dy=du when dy=du, that indicates no real need for substitution. if you know what an integral of 1/y is, then you should be also able to tell me what an integral of 1/(y+1) and the integral of 1/(y-1) is

OpenStudy (anonymous):

i need help with that, all i can think of is 1/(2y+y) and 1/(2y-2)

OpenStudy (solomonzelman):

lets take it bit by bit.

OpenStudy (solomonzelman):

What is an integral of 1/y ?

OpenStudy (anonymous):

y^2/2?

OpenStudy (solomonzelman):

that is integral of y, but what is integral of 1/y ?

OpenStudy (anonymous):

1/2y^2

OpenStudy (anonymous):

2/2y^2 actually

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \int \frac{1}{y} ~dy~~=\ln \left| y\right| \color{gray}{+C}}\)

OpenStudy (solomonzelman):

and, thus \(\large\color{black}{ \displaystyle \int \frac{1}{y\pm 1} ~dy~~=\ln \left| y\pm 1\right| \color{gray}{+C}}\)

OpenStudy (anonymous):

ok so its ln y

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \int \frac{1}{4(y+1)} +\frac{3}{4(y-3)} dy~~=\frac{1}{4}\ln \left| y+1\right|+\frac{3}{4}\ln \left| y-3\right| \color{gray}{+C}}\)

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \displaystyle \int_{4}^{8} \frac{1}{4(y+1)} +\frac{3}{4(y-3)} dy~~=\left(~\frac{1}{4}\ln \left| y+1\right|+\frac{3}{4}\ln \left| y-3\right| ~\right){{\huge |}_{4}^{8}}}\)

OpenStudy (solomonzelman):

now plug in the limits

OpenStudy (anonymous):

ok thank you for all you help

OpenStudy (solomonzelman):

you welcome!

OpenStudy (anonymous):

ok i need help

OpenStudy (anonymous):

nvm